Tyler received $1200 from his parents as a graduation present. He invested part of it at 4% interest, and he invested the remainder at 6%. If the total yearly interest amounted to $62, how much did he invest at each rate?
Answer by mananth(16946)

Part I 4.00% per annum ------------- Amount invested =x
Part II 6.00% per annum ------------ Amount invested = y
1200
Interest----- 62

Part I 4.00% per annum ---x
Part II 6.00% per annum ---y
Total investment
x + 1 y= 1200 -------------1
Interest on both investments
4.00% x + 6.00% y= 62
Multiply by 100
4 x + 6 y= 6200.00 --------2
Multiply (1) by -4
we get
-4 x -4 y= -4800.00
Add this to (2)
0 x 2 y= 1400
divide by 2
y = 700
Part I 4.00% $ 500
Part II 6.00% $ 700

CHECK
500 --------- 4.00% ------- 20.00
700 ------------- 6.00% ------- 42.00
Total -------------------- 62.00

m.ananth@hotmail.ca